OpenStudy (anonymous):

Let tan β = 4 Find cot β. A.4/17 (√17) B. None of these C. 4 D.1/4 E.√17/4 F.√17/17

zepdrix (zepdrix):

All the way down to F? Wow they sure give you a lot of choices :)

OpenStudy (anonymous):

\[\cot (\theta) = \frac{ 1 }{ \tan (\theta) }\]

OpenStudy (anonymous):

\[\cot(\theta) = \frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

cot is the inverse function of tan

OpenStudy (anonymous):

Anyways, hope that helps